OpenStudy (anonymous):

19. A box contains 95 pink rubber bands and 90 brown rubber bands. You select a rubber band at random from the box. Find each probability. Write the probability as a fraction in simplest form. a. Find the theoretical probability of selecting a pink rubber band. b. Find the theoretical probability of selecting a brown rubber band. c. You repeatedly choose a rubber band from the box, record the color, and put the rubber band back in the box. The results are shown in the table below. Find the experimental probability of each color based on the table.

OpenStudy (anonymous):

Table: OUTCOME OCCURANCES Pink 36 Brown 33

OpenStudy (anonymous):

Can you show me how to solve this by drawing

OpenStudy (kropot72):

a. Add up to find the total number of number bands of both colors in the box. The theoretical probability of selecting a pink rubber band is given by: \[\frac{number\ of\ pink\ bands}{total\ number\ of\ bands}\]

OpenStudy (anonymous):

36/69=.52

OpenStudy (anonymous):

33/69=.47

OpenStudy (kropot72):

b. The theoretical probability of selecting a brown rubber band is given by: \[\frac{number\ of\ brown\ bands}{total\ number\ of\ bands}\]

OpenStudy (anonymous):

c: thanks. And now the next step?

OpenStudy (kropot72):

Your answer for c. 'the experimental probability of pink' is correct. However the experimental probability of brown is 33/69 = 0.48.

OpenStudy (anonymous):

OH yeah you would round because of the 8

OpenStudy (kropot72):

The ways to solve a. and b. are given in my previous posts.

OpenStudy (anonymous):

So, a would be .52 and b would be .48

OpenStudy (anonymous):

I found the theoretical probabilities on how to find it, now how would I do experimental?

OpenStudy (kropot72):

I get 0.51 for a. and 0.49 for b.

OpenStudy (anonymous):

Can you maybe show me how you got it? o: I know I'm such a pain XD

OpenStudy (kropot72):

You have the theoretical probabilities for a. and b. and the experimental probabilities for c. That is all the question asked for.

OpenStudy (kropot72):

a. 95/185 = 0.51351 = 0.51 (rounded to 2 decimal places) b. 90/185 = 0.49189 = 0.49 ( rounded to 2 decimal places)

OpenStudy (anonymous):

:D:D:D:D:D:D:D:D Thanks!!!!

OpenStudy (kropot72):

You're welcome :)